40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8568     Entry time: Tue May 14 01:13:35 2013
Author: Jenne 
Type: Update 
Category: 40m Upgrading 
Subject: TRY realigned 

Koji noticed that earlier this afternoon the Yarm ASS was working, but then after dinner it was no longer working.  I saw that the ETMY trans camera beam was clipped.  These things precipitated a visit to the Yend station. 

I saw that the beam on the optic that steers the camera beam to the camera was very, very low, almost falling off the optic.  The only mirror which steers to this optic is the harmonic separator which reflects the IR, and transmits the green.  I turned the pitch knobs on the harmonic separator until the beam was roughly centered on all 3 optics between the separator and the camera (BS to QPD, BS to TRYDC and Y1 for camera).  The yaw was fine, so I didn't touch it.

I then adjusted the steering mirror to the camera, and the BS pointing to the DC PD.  I have not touched the BS pointing to the QPD.  Once the beam was on the TRY PD, Koji ran the ASS script, and I recentered the beam on the DC PD.  During this time, Koji had the Yarm triggering using -1 in the POYDC element of the matrix.

The harmonic separator is not mounted in a nice way (I'm assuming that Annalisa is in the middle of things, and she'll get back to it after the green work), so the TRY PD and camera will need to be aligned again, so I didn't do any ASS-recentering-ASS iteration tonight.

The Yarm ASS works nicely again, getting TRY to ~0.89 . 

ELOG V3.1.3-